OpenStudy (anonymous):

Whats The Rate Of Change For 5x + 3y = -2 How Did You Find it?

OpenStudy (anonymous):

Please explain as much as you can :)

OpenStudy (anonymous):

isolate y 3y = -5x - 2 y = -5/3 x - 2/3 slope is -5/3 since the equation is y = mx + b, where m is the slope

OpenStudy (anonymous):

so the answer is y=-5/3x-2?

OpenStudy (anonymous):

answre is -5/3 since you are only looking for the rate of change